Jordan Inc has the following balance sheet and income statement data: Cash $ 14,000 Accounts payable $ 42,000 Receivables 70,000 Other current liab. 28,000 Inventories 280,000 Total CL $ 70,000 Total CA $364,000 Long-term debt 140,000 Net fixed assets 126,000 Common equity 280,000 Total assets $490,000 Total liab. and equity $490,000 Sales $280,000 Net income $ 21,000 ​ The new CFO thinks that inventories are excessive and could be lowered sufficiently to cause the current ratio to equal the industry average, 2.75, without affecting either sales or net income. Assuming that inventories are sold off and not replaced to get the current ratio to the target level, and that the funds generated are used to buy back common stock at book value, by how much would the ROE change?

a. 11.26%

b. 13.07%

c. 11.85%

d. 12.45%

e. 13.72%
Answer #1

Total current assets = 364,000

Total current liabilities = 70,000

so present current ratio = 364,000 / 70000 = 5.2 (current ratio = current assets / current liabilities)

targeted current ratio = 2.75

so required current assets = 70,000 x 2.75 = $192,500

reduction in current assets = 364,000 - 192,500 = 171,500

reduction in inventory = 171,500 (inventories are lowered)

also reduction in common equity is reduction in inventory

new common equity after buy back = 280,000 - 171500

= 108,500

ROE = net income / equity

ROE before buy back = 21,000 / 280,000 = 7.5%

ROE after buy back = 21,000 / 108,500 = 19.35%

change in ROE = 19.35 - 7.5 = 11.85%

Option c is correct answer